# Towards Federated Learning with on-device Training and Communication in 8-bit Floating Point

### v1

@@ 2. Method
$\mathbf{w}_{1},\alpha_{1},\beta_{1},Q_{\text{det}},Q_{\text{rand}}$
for $t=1,...,T$ do
     Sample a set $\mathcal{P}_{t}\in[K]$ of $P$ active devices
     for each client $k\in\mathcal{P}_{t}$ do
         Receive $Q_{\text{rand}}(\mathbf{w}_{t};\alpha_{t}),\alpha_{t},\beta_{t}$ from server
         $\{\mathbf{w}^{k}_{t+1},\alpha^{k}_{t+1},\beta^{k}_{t+1}\}\leftarrow\text{LocalUpdate}(\mathbf{w}_{t},Q_{\text{det}};\alpha_{t},\beta_{t},\mathcal{D}_{k})$
         Send $Q_{\text{rand}}(\mathbf{w}^{k}_{t+1};\alpha^{k}_{t+1}),\alpha^{k}_{t+1},\beta^{t+1}_{i}$ to server
     end for
     Compute $m_{t}=\sum_{k\in\mathcal{P}_{t}}n_{k}$, $\beta_{t+1}\leftarrow\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\beta^{k}_{t+1}$
      Compute $\mathbf{w}_{t+1}\leftarrow\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}Q_{\text{rand}}(\mathbf{w}^{k}_{t+1};\alpha^{k}_{t+1})$
      $\alpha_{t+1}\leftarrow\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\alpha^{k}_{t+1}$
     $\{\mathbf{w}_{t+1},\alpha_{t+1}\}\leftarrow\textsc{ServerOptimize}(\{\alpha^{k}_{t+1},\mathbf{w}^{k}_{t+1}\}_{k\in\mathcal{P}_{t}})$
end for
Evaluate on $\mathbf{w}_{T+1},\alpha_{T+1},\beta_{T+1}$
Server-Side Optimization (ServerOptimize). The standard federated average of the weights in the un-quantized scenario corresponds to the minimization of weighted mean squared error (MSE) between the server parameters and the client parameters. However, when the server parameters are quantized before transmission to the clients, this property no longer holds. We therefore propose a modification to the server-side model aggregation, where the MSE is explicitly minimized. This can be done without communication overhead since all computations are done on the server. We are leveraging the computing power of the server to do more optimization since the server typically possesses more computing power than a device. At time step $t$, we perform model/parameters aggregation to obtain $\mathbf{w}_{t+1},\alpha_{t+1}$ for the next communication round by minimizing the following mean-squared error (MSE).
Note that when there is no communication quantization, the closed-form solution to ServerOptimize is the federated average $\mathbf{w}_{t+1}=\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\mathbf{w}_{t}^{k}$. Since the problem above has no closed-form solution for the quantized communication case, we use the alternating minimization approach to optimize $\mathbf{w}$ and $\alpha$: First, we optimize the model weights $\mathbf{w}$ by minimizing (4) using a fixed number of gradient descent steps, while fixing $\alpha$ to $\alpha_{t+1}=\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\alpha_{t+1}^{k}$.
(4) $\mathbf{w}_{t+1}=\operatorname*{arg\,min}_{\mathbf{w}}\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\|Q_{\text{rand}}(\mathbf{w};\alpha_{t+1})-Q_{\text{rand}}(\mathbf{w}^{k}_{t};\alpha^{k}_{t})\|^{2}_{2}.$
(5) $\alpha_{t+1}=\operatorname*{arg\,min}_{\alpha\in S}\sum_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\|Q_{\text{rand}}(\mathbf{w}_{t};\alpha)-Q_{\text{rand}}(\mathbf{w}^{k}_{t};\alpha^{k}_{t})\|^{2}_{2}.$
Overall algorithm. A summary of our proposed FP8 FedAvg with unbiased communication (FP8FedAvg-UQ) method is presented in Algorithm 1, where the optional server-optimization step (UQ+) corresponds to replacing the standard federated averaging of weight and range parameters with our two-step MSE minimization optimization in equations (4) and (5). It is important to note that our method involves two different quantization operators. On-device QAT uses a deterministic and biased quantizer while the communication part adopts its stochastic counterpart which is unbiased. In the next section, we will give a convergence analysis result for FP8FedAvg-UQ and show that these design choices are well-motivated from a theory point of view.
3. Convergence analysis and theoretical motivations
We briefly state our main convergence theorem here and refer to the Appendix for the formal assumption definitions and proof. Please note that we make the simplifying assumption to only consider weight quantization in our proof, which is standard for this type of theoretical analysis.
@@ 3. Convergence analysis and theoretical motivations
For convex and $L$-smooth federated losses in (1) with $G$-bounded unbiased stochastic gradients using an FP8 deterministic quantization method during training and an FP8 unbiased quantization method with bounded scales for model communication, the objective gap $E\left[F(Q_{\text{rand}}(\mathbf{w}_{\tau}))-F(\mathbf{w}_{*})\right]$ decreases at a rate of
$O\biggl{(}\underbrace{\frac{\|\mathbf{w}_{1}-\mathbf{w}_{*}\|^{2}_{2}}{\sqrt{TU}}+\frac{G^{2}\sqrt{U}}{\sqrt{T}}+\frac{UG^{2}L}{T}}_{\mathcal{T}_{1}}+\underbrace{\frac{GU^{2.5}S\sqrt{d}L}{\sqrt{T}}}_{\mathcal{T}_{2}}+\underbrace{S\sqrt{d}G}_{\mathcal{T}_{3}}\biggl{)}$
where $\tau$ is uniformly sampled from $\{1,2,\dotsc,T\}$, $T$ is the number of rounds, $U$ is the total number of updates done in each round, the quantization scales $s_{i}$ are uniformly bounded by $S$, $w_{1}$ is the initial model, and $w_{*}$ is an optimal solution of (1).
@@
$\mathcal{T}_{1}$ is a term similar to SGD convergence where it decreases with $O\left(\frac{1}{\sqrt{T}}\right)$ and depends on the bound on the second moment of stochastic gradient $G$, smoothness $L$, as well as the $\ell_{2}$-distance between the initial model $w_{1}$ and the optimal solution $w_{*}$.
@@
$\mathcal{T}_{2}$ and $\mathcal{T}_{3}$ are due to quantization. Due to (2) and the definition of $S$, the terms $\mathcal{T}_{2}$ and $\mathcal{T}_{3}$ exponentially decay when the number of mantissa bits $m$ increases, i.e., $\mathcal{T}_{2}\propto 2^{-m},\mathcal{T}_{3}\propto 2^{-m}$.
@@
We emphasize that unbiased quantization during communication is crucial. In the case of biased communication, the convergence proof does not hold and one can construct even diverging cases (Beznosikov et al., 2023) for FedAvg. To ensure convergence for biased communication, we need more sophisticated techniques such as error feedback (Richtárik et al., 2021).
@@
Deterministic quantization is used during training. We bound the norm of QAT quantization error in the proof. Since deterministic quantization has a smaller error norm than stochastic one, we use deterministic quantization during training.
@@
4. Experiments and Ablation Studies
@@
Setup. We test our method on two different tasks: image classification on CIFAR10 and CIFAR100 (Krizhevsky, 2009) and keyword spotting on Google SpeechCommands v2 (Warden, 2018) (35-word task). For image classification, we use the LeNet and ResNet18 (He et al., 2016) architectures, and for keyword spotting we use MatchboxNet3x1x64 (Majumdar and Ginsburg, 2020) and the Transformer-based KWT-1 model (Berg et al., 2021). Furthermore, we replace batch normalization layers in the convolutional networks with GroupNorm (Wu and He, 2018), since this is known to work better in a federated setting with skewed data distributions (Hsieh et al., 2020). For all FP8 experiments, we use one signed bit, $m=3$ and $e=4$ bits for the mantissa and exponent respectively.
@@
For each dataset, we evaluate our method in both i.i.d. and non-i.i.d. distributions of the datasets across clients. In the i.i.d scenarios, we use $K=100$ clients, a participation rate of $C=0.1$ in each round and train for $R=1000$ rounds with a local batch size of $B=50$, where each client trains for 5 local epochs. In the non-i.i.d. image classification setting we sample the local datasets from a Dirichlet distribution with a concentration parameter of 0.3.
@@
For the keyword spotting task, we use a similar setup as (Li et al., 2022) and split the dataset based on the speaker-id of each utterance in order to obtain a more realistic heterogenity across local datasets. This results in a total of $K=2112$ clients for the 35-word task. In order to get a similar total training steps as in the i.i.d scenarion, we reduce the participation rate to $C=0.01$, the number of rounds to $R=500$ and the number of local gradient updates to 50.
When training models on image classification, we use SGD as the local optimizer with a constant learning rate of 0.1, weight decay of 0.001 and random cropping and horizontal flipping for data augmentation. On SpeechCommands, we adhere to the training setup used in (Berg et al., 2021), with AdamW (Loshchilov and Hutter, 2018) as local optimizer with an initial learning rate of 0.001 and a cosine decay scheduler, weight decay of 0.1, and apply SpecAugment (Park et al., 2019) to the mel-frequency cepstrum coefficients.
When applying client-side FP8 QAT, we quantize all parameters in the network except biases and parameters in normalization layers, e.g. GroupNorm and LayerNorm. This results in a negligible impact on the client-server communication, since these parameters account for ¡ 2 % of the total parameter count in the models. In addition, when performing server-side optimization of weight aggregation, we use 5 gradient descent steps when solving (4), where the learning rate was selected using grid search in $\{0.01,0.1,1\}$, and 50 grid points when solving (5).
@@
Table 1. Final test accuracy and communication gain compared to FP32 FedAvg for our proposed methods on CIFAR10/CIFAR100 and Google SpeechCommands.
@@
LeNet i.i.d. $82.1\pm 0.1$ / 1×1$\times$1 × $82.0\pm 0.1$ / 4.1×4.1$\times$4.1 × $82.2\pm 0.3$ / 4.7×4.7$\times$4.7 ×
Dir(0.3) $77.1\pm 0.4$ / 1×1$\times$1 × $77.3\pm 0.9$ / 3.9×3.9$\times$3.9 × $77.7\pm 0.5$ / 3.9×3.9$\times$3.9 ×
ResNet18 i.i.d $92.0\pm 0.1$ / 1×1$\times$1 × $91.1\pm 0.2$ / 3.4×3.4$\times$3.4 × $92.0\pm 0.1$ / 3.9×3.9$\times$3.9 ×
Dir(0.3) $85.5\pm 0.5$ / 1×1$\times$1 × $87.4\pm 0.7$ / 5.2×5.2$\times$5.2 × $87.5\pm 0.5$ / 5.2×5.2$\times$5.2 ×
@@
LeNet i.i.d. $43.0\pm 0.3$ / 1×1$\times$1 × $44.8\pm 0.4$ / 6.0×6.0$\times$6.0 × $44.9\pm 0.5$ / 6.0×6.0$\times$6.0 ×
Dir(0.3) $38.3\pm 0.7$ / 1×1$\times$1 × $41.1\pm 0.3$ / 9.1×9.1$\times$9.1 × $41.3\pm 0.7$ / 9.5×9.5$\times$9.5 ×
ResNet18 i.i.d $64.6\pm 0.3$ / 1×1$\times$1 × $64.0\pm 0.2$ / 3.5×3.5$\times$3.5 × $64.6\pm 0.1$ / 4.0×4.0$\times$4.0 ×
Dir(0.3) $56.1\pm 0.7$ / 1×1$\times$1 × $55.4\pm 0.6$ / 3.6×3.6$\times$3.6 × $55.5\pm 0.6$ / 3.6×3.6$\times$3.6 ×
@@
MatchboxNet i.i.d. $91.5\pm 0.3$ / 1×1$\times$1 × $90.0\pm 0.4$ / 3.5×3.5$\times$3.5 × $90.8\pm 0.4$ / 3.4×3.4$\times$3.4 ×
speaker-id $79.6\pm 0.7$ / 1×1$\times$1 × $75.4\pm 0.6$ / 3.1×3.1$\times$3.1 × $77.0\pm 1.3$ / $\times3$
KWT-1 i.i.d $91.4\pm 0.4$ / 1×1$\times$1 × $89.2\pm 0.3$ / 2.3×2.3$\times$2.3 × $90.7\pm 0.2$ / 2.9×2.9$\times$2.9 ×
speaker-id $83.2\pm 0.2$ / 1×1$\times$1 × $79.6\pm 0.5$ / 2.9×2.9$\times$2.9 × $82.4\pm 0.8$ / 3.7×3.7$\times$3.7 ×
Results. The results are shown in Table 1, where we present the final centralized test accuracy as well as the communication gain for each model and dataset, with and without server-side optimization, across three random seeds. In order to compare communication costs, we do not pick a common accuracy threshold for all methods, but instead calculate the gains individually for each method as the reduction in communicated bytes compared to FP32 training at the maximum accuracy reached by both FP32 and FP8. In Table 1 it can be seen that for most datasets and methods, FP8-FedAvg-UQ achieves similar test accuracy as the FP32 baseline, which results in communication gains around 4x on average. Note that even though FP8 quantization sometimes results in a small accuracy drop, a large communication gain is still possible due to less data being transferred in each communication round. However, in certain cases, for example when applying FP8 quantization to LeNet on CIFAR100, accuracy increases significantly compared to the FP32 baseline. For these experiments, we observed that FP8 quantization to some extent prevented overfitting to the local client datasets, and thereby acted as a regularizer. This effect of quantization has been observed in other studies as well (AskariHemmat et al., 2022). Finally, we note that the server-side optimization in most scenarios yields additional performance improvements, which results in communication gains greater than 2.9x compared to FP32 across all tasks and models.
Ablation studies. Next, we ablate the use of deterministic and stochastic quantization in order to validate our design choices. Table 2 shows the server test accuracy when using the two different quantization methods for the on-device QAT training. Deterministic quantization works best here, which can also be understood intuitively, since in each forward pass through the network, deterministic quantization results in a smaller quantization error. We refer to Appendix B for more details about QAT convergence.
We also ablate the effect of deterministic and stochastic quantization in server-device communication, and it is clear that stochastic quantization results in both higher accuracy and gain. This is in agreement with Remark 3, which shows that stochastic quantization is important for the convergence of our algorithm. In addition, we show the server test accuracy as a function of communication cost for different methods in Figure 2. Here we can clearly see the gain arising from quantized communication, as well as the benefits of stochastic quantization and server-side optimization.

### v2

@@ 3.6 Overall algorithm
[…]
1:for $t=1,...,T$ do
2:  Sample a set $\mathcal{P}_{t}\in[K]$ of $P$ active devices
3:  for each client $k\in\mathcal{P}_{t}$ do
4:   Receive $Q_{\text{rand}}(\mathbf{w}_{t};\alpha_{t}),\alpha_{t},\beta_{t}$ from server
5:   $\{\mathbf{w}^{k}_{t+1},\alpha^{k}_{t+1},\beta^{k}_{t+1}\}\leftarrow\textsc{LocalUpdate}(\mathbf{w}_{t},Q_{\text{det}};\alpha_{t},\beta_{t},\mathcal{D}_{k})$
6:   Send $Q_{\text{rand}}(\mathbf{w}^{k}_{t+1};\alpha^{k}_{t+1}),\alpha^{k}_{t+1},\beta^{k}_{t+1}$ to server
7:  end for
8:  Compute $m_{t}=\sum\limits_{k\in\mathcal{P}_{t}}n_{k}$
9:  Compute $\beta_{t+1}\leftarrow\sum\limits_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\beta^{k}_{t+1}$
10:   Compute $\mathbf{w}_{t+1}\leftarrow\sum\limits_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}Q_{\text{rand}}(\mathbf{w}^{k}_{t+1};\alpha^{k}_{t+1})$
11:   Compute $\alpha_{t+1}\leftarrow\sum\limits_{k\in\mathcal{P}_{t}}\frac{n_{k}}{m_{t}}\alpha^{k}_{t+1}$
12:  $\{\mathbf{w}_{t+1},\alpha_{t+1}\}\leftarrow\textsc{ServerOptimize}(\{\mathbf{w}^{k}_{t+1},\alpha^{k}_{t+1},n_{k}\}_{k\in\mathcal{P}_{t}},Q_{\text{rand}})$
13:end for
14: Evaluate on $\mathbf{w}_{T+1},\alpha_{T+1},\beta_{T+1}$
4 Convergence analysis and theoretical motivations
We briefly state our main convergence theorem here and refer to the Appendix for formal assumption definitions and proof. Please note that we make the simplifying assumption to only consider weight quantization in our proof, which is standard for this type of theoretical analysis.
@@ 4 Convergence analysis and theoretical motivations
For convex and $L$-smooth federated losses in (1) with $G$-bounded unbiased stochastic gradients using an FP8 deterministic quantization method during training and an FP8 unbiased quantization method with bounded scales for model communication, the objective gap $\mathbb{E}\left[F(Q_{\text{rand}}(\mathbf{w}_{\tau}))-F(\mathbf{w}_{*})\right]$ decreases at a rate of
$O\biggl{(}\underbrace{\frac{\Delta^{2}+G^{2}U}{\sqrt{TU}}+\frac{UG^{2}L}{T}}_{\mathcal{T}_{1}}+\underbrace{\frac{GU^{2.5}S\sqrt{d}L}{\sqrt{T}}}_{\mathcal{T}_{2}}+\underbrace{S\sqrt{d}G}_{\mathcal{T}_{3}}\biggl{)},$
where $\Delta=\|\mathbf{w}_{1}-\mathbf{w}_{*}\|_{2}$ is the initial and optimal model difference, $\tau$ is uniformly sampled from $\{1,2,\dotsc,T\}$, $T$ is the number of rounds, $U$ is the total number of updates done in each round, the quantization scales $s_{i}$ are uniformly bounded by $S$, $\mathbf{w}_{1}$ is the initial model, and $\mathbf{w}_{*}$ is an optimal solution of (1).
Proof Structure. The proof builds on upper bounding a drift quantity similar to the one defined in karimireddy2020scaffold as [^2] See Appendix C.2 on ${\mathbf{w}}_{t,u}^{k}$ definition for inactive devices.,
$V_{t}=\frac{1}{KU}\sum_{k\in[K]}\sum_{u\in[U]}E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{w}}_{t,u}^{k}\right)\right\|^{2}_{2}.$
Note that if local models diverge, we get a higher $V_{t}$. If there is no quantization and local models converge, we get $V_{t}=0$. We focus on $V_{t}$ and the server model $\mathbf{w}$ in a single communication round and give the following Lemma as,
Lemma 1.
For a setting that satisfies the assumptions listed in Theorem 4.1, we have,
$V_{t}\leq 18$ $U^{3}S\sqrt{d}G\eta_{t}+9U^{2}\eta_{t}^{2}G^{2},$ (9)
$E\|\mathbf{w}_{t+1}-\mathbf{w}_{*}\|^{2}_{2}-E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}\leq$ $\eta_{t}LUV_{t}+2S\sqrt{d}GU\eta_{t}+\eta_{t}^{2}U^{2}G^{2}$
$-2U\eta_{t}E\left(F\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F\left(\mathbf{w}_{*}\right)\right),$ (10)
where $\eta_{t}$ is the learning rate. By combining the equations with proper coefficients, we get a telescoping sum on the difference between the server model and the optimal model, as well as the accumulation of the loss with respect to the optimal model. This leads to Theorem 4.1 and we provide the full proof in Appendix C.
@@
$\mathcal{T}_{1}$ is a term similar to SGD convergence where it decreases with $O\left(\frac{1}{\sqrt{T}}\right)$ and depends on the bound on the second moment of stochastic gradient $G$, the smoothness $L$, as well as the $\ell_{2}$-distance between the initial model $w_{1}$ and the optimal solution $w_{*}$.
@@
$\mathcal{T}_{2}$ and $\mathcal{T}_{3}$ are terms that arise due to quantization. Due to (4) and the definition of $S$, the terms $\mathcal{T}_{2}$ and $\mathcal{T}_{3}$ exponentially decay when the number of mantissa bits $m$ increases, i.e., $\mathcal{T}_{2}\propto 2^{-m},\mathcal{T}_{3}\propto 2^{-m}$.
@@
We emphasize that unbiased quantization during communication is crucial. In the case of biased communication, the convergence proof does not hold and one can construct even diverging cases beznosikov2023biased for FedAvg. To ensure convergence for biased communication, we need more sophisticated techniques such as error feedback richtarik2021ef21 .
@@
Deterministic quantization is used during training. We bound the norm of QAT quantization error in the proof. Since deterministic quantization has a smaller error norm than stochastic one, this motivates us to use deterministic quantization during training.
@@
5 Experiments and Ablation Studies
We test our proposed method on two different tasks: image classification and keyword spotting. For each task, we use two different models and perform experiment using both an i.i.d. and a non-i.i.d. dataset split across clients. In the i.i.d. setup, the dataset is split randomly across the set of clients. In the non-i.i.d. setup, we simulate a more realistic heterogenity across clients which is specific to each dataset.
5.1 Datasets and models
@@
Image classification. For image classification, we use the CIFAR10 and CIFAR100 datasets Krizhevsky09learningmultiple , which consist of 60 000 examples of 32x32 color images divided into 10 and 100 different classes respectively. For this task, we adopt two different convolutional networks: 1) LeNet lecun1998gradient , which has 800K parameters, and ResNet18 he2016deep , with 11M parameters. For the ResNet model, we replace batch normalization after the convolutional layers with GroupNorm wu2018group , since this is known to work better in a federated setting with skewed data distributions hsieh2020non .
Figure 3: Illustration of the QAT method for a linear/convolutional layer. FP32 master weights are quantized using the range parameters $\alpha$, before being applied to the already quantized input. We then apply addition of bias, an optional normalization layer and the activation function in FP32, before quantizing to FP8 again using the range parameters $\beta$. In a real application, these operations could be performed using mixed-precision hardware.
@@
In the (independent and identical and distributed) i.i.d. scenario, we use $K=100$ clients, a participation rate of $C=0.1$ in each round and train for $R=1000$ rounds with a local batch size of $B=50$, where each client trains for 5 local epochs. In the non-i.i.d. image classification setting we sample the local datasets from a Dirichlet distribution with a concentration parameter of 0.3. Furthermore, we use SGD as the local optimizer with a constant learning rate of 0.1, weight decay of 0.001 and random cropping and horizontal flipping for data augmentation.
Keyword spotting. In order to apply our method to a more realistic federated learning task and more advanced model architectures, we resort to keyword spotting, where the task is to classify short snippets of audio as words in a small dictionary. For this, we use the Google SpeechCommands v2 dataset, which consists of 105 000 one-second recordings belonging to one of 35 classes warden2018speech . Examples of classes in the dictionary include short words like “go”, “yes” and “on”.
For the keyword spotting task, we train two different models: MatchboxNet3x1x64 majumdar20_interspeech , which uses 1D time-channel separable convolutions with skip-connections, as well as the Keyword Transformer (KWT-1) berg21_interspeech model, which consists of sequential transformer layers with time-domain self-attention. These two models have 350K and 450K parameters respectively and they both use mel-frequency cepstral coefficients (MFCCs) as input features.
A known problem when training transformers is that SGD optimization requires many iterations and may fail to converge to a good solution kunstner2023noise . Due to the relatively small number of local updates in a training round, SGD is therefore not a suitable optimization method for transformers in this setup. To achieve better training convergence, we instead use the momentum-based AdamW loshchilov2018decoupled as local optimizer for the clients, where the states of the local optimizer for each active client are reset at the start of each communication round. Furthermore, we use an initial learning rate of 0.001 with a cosine decay scheduler and a weight decay of 0.1. For data augmentation, we apply SpecAugment park19e_interspeech to the MFCCs during training.
@@
When training in the i.i.d. scenario, we use the same setup for keyword spotting as for image classification. However, since the SpeechCommands datasets provides speaker identity as well, we can simulate a more realistic heterogenity for the non-i.i.d. dataset split. We therefore use the setup proposed in li22g_interspeech , such that all recordings from a given speaker are assigned to a single client. This results in a total of $K=2112$ clients. In order to get a similar number of total training steps as in the i.i.d scenario, we reduce the participation rate to $C=0.01$, the number of rounds to $R=500$ and the number of local gradient updates to 50.
5.2 Mixed Precision Quantization Implementation
Since most of the bandwidth during the communication phase between the server and clients is used for transmitting the model weights in the fully connected and convolutional layers, we focus on learning FP8 representations for these. Hence, we simulate a mixed-precision training scenario where these parameters are trained in FP8, whereas bias parameters and normalization layer parameters (GroupNorm for convolutional networks and LayerNorm for transformers) are trained in FP32, since initial experiments showed these to be sensitive to quantization. Note that this only results in a negligible impact on the client-server communication, since these parameters account for less than 2 % of the total parameter count in the models.
An illustration of how mixed-precision training can be simulated using QAT is shown in Figure 3. The diagram illustrates the necessary operations for quantizing most layers in the networks we use. A notable exception is the self-attention layers in the transformer model. In these layers, we perform calculations of the keys, queries and values, as well as calculation of the dot-product attention scores in FP8. Softmax normalization of the attention scores is done in FP32, similar to how it is done for other activation functions.
[…]
5.3 Results
The results are shown in Table 1, where we present the final centralized test accuracy and standard deviation across three random seeds, as well as the average communication gain compared to an FP32 baseline. In order to compare communication costs, we do not pick a common accuracy threshold for all methods, but instead calculate the gains individually for each method as the reduction in communicated bytes compared to FP32 training at the maximum accuracy reached by both FP32 and FP8.
In Table 1 it can be seen that for most datasets and methods, FP8-FedAvg-UQ achieves similar test accuracy as the FP32 baseline, which results in communication gains of 4.2x on average, and for some experiments larger than 9x. Note that even though FP8 quantization sometimes results in a small accuracy drop, a large communication gain is still possible due to less data being transferred in each communication round. However, in certain cases, for example when applying FP8 quantization to LeNet on CIFAR100, accuracy increases significantly compared to the FP32 baseline. For these experiments, we observed that FP8 quantization to some extent prevented overfitting to the local client datasets, and thereby acted as a regularizer. This effect of quantization has been observed in other studies as well askarihemmat2022qreg .
Table 1 further shows results for i.i.d. and a more realistic non-i.i.d. settings. Here, we see an expected accuracy drop when moving from the i.i.d. to the non-i.i.d. setting, yet our method shows gains in both scenarios. We note that our method could potentially be combined with more advanced federated learning algorithms in heterogenous settings to improve accuracy levels in the non-i.i.d. setting as well.
Finally, we note that the server-side optimization in most scenarios yields additional performance improvements, with an average gain of 4.5x. The impact of this additional optimization is most notable when the FP8 quantization results in an accuracy drop compared to FP32. On the contrary, when there is no accuracy loss, the improvement due to server optimization is smaller. Overall, the quantized communication in combination with server-side optimization results in communication gains greater than 2.9x compared to FP32 across all tasks and models.
@@ 5.3 Results
Table 1: Final test accuracy and communication gain compared to FP32 FedAvg for our proposed methods on CIFAR10/CIFAR100 and Google SpeechCommands.
@@
LeNet i.i.d. 82.1±0.182.1$\pm$ 0.182.1 ± 0.1 / 1×1$\times$1 × 82.0±0.182.0$\pm$ 0.182.0 ± 0.1 / 4.1×4.1$\times$4.1 × 82.2±0.382.2$\pm$ 0.382.2 ± 0.3 / 4.7×4.7$\times$4.7 ×
Dir(0.3) 77.1±0.477.1$\pm$ 0.477.1 ± 0.4 / 1×1$\times$1 × 77.3±0.977.3$\pm$ 0.977.3 ± 0.9 / 3.9×3.9$\times$3.9 × 77.7±0.577.7$\pm$ 0.577.7 ± 0.5 / 3.9×3.9$\times$3.9 ×
ResNet18 i.i.d 92.0±0.192.0$\pm$ 0.192.0 ± 0.1 / 1×1$\times$1 × 91.1±0.291.1$\pm$ 0.291.1 ± 0.2 / 3.4×3.4$\times$3.4 × 92.0±0.192.0$\pm$ 0.192.0 ± 0.1 / 3.9×3.9$\times$3.9 ×
Dir(0.3) 85.5±0.585.5$\pm$ 0.585.5 ± 0.5 / 1×1$\times$1 × 87.4±0.787.4$\pm$ 0.787.4 ± 0.7 / 5.2×5.2$\times$5.2 × 87.5±0.587.5$\pm$ 0.587.5 ± 0.5 / 5.2×5.2$\times$5.2 ×
@@
LeNet i.i.d. 43.0±0.343.0$\pm$ 0.343.0 ± 0.3 / 1×1$\times$1 × 44.8±0.444.8$\pm$ 0.444.8 ± 0.4 / 6.0×6.0$\times$6.0 × 44.9±0.544.9$\pm$ 0.544.9 ± 0.5 / 6.0×6.0$\times$6.0 ×
Dir(0.3) 38.3±0.738.3$\pm$ 0.738.3 ± 0.7 / 1×1$\times$1 × 41.1±0.341.1$\pm$ 0.341.1 ± 0.3 / 9.1×9.1$\times$9.1 × 41.3±0.741.3$\pm$ 0.741.3 ± 0.7 / 9.5×9.5$\times$9.5 ×
ResNet18 i.i.d 64.6±0.364.6$\pm$ 0.364.6 ± 0.3 / 1×1$\times$1 × 64.0±0.264.0$\pm$ 0.264.0 ± 0.2 / 3.5×3.5$\times$3.5 × 64.6±0.164.6$\pm$ 0.164.6 ± 0.1 / 4.0×4.0$\times$4.0 ×
Dir(0.3) 56.1±0.756.1$\pm$ 0.756.1 ± 0.7 / 1×1$\times$1 × 55.4±0.655.4$\pm$ 0.655.4 ± 0.6 / 3.6×3.6$\times$3.6 × 55.5±0.655.5$\pm$ 0.655.5 ± 0.6 / 3.6×3.6$\times$3.6 ×
@@
MatchboxNet i.i.d. 91.5±0.391.5$\pm$ 0.391.5 ± 0.3 / 1×1$\times$1 × 90.0±0.490.0$\pm$ 0.490.0 ± 0.4 / 3.5×3.5$\times$3.5 × 90.8±0.490.8$\pm$ 0.490.8 ± 0.4 / 3.4×3.4$\times$3.4 ×
speaker-id 79.6±0.779.6$\pm$ 0.779.6 ± 0.7 / 1×1$\times$1 × 75.4±0.675.4$\pm$ 0.675.4 ± 0.6 / 3.1×3.1$\times$3.1 × 77.0±1.377.0$\pm$ 1.377.0 ± 1.3 / $\times3$
KWT-1 i.i.d 91.4±0.491.4$\pm$ 0.491.4 ± 0.4 / 1×1$\times$1 × 89.2±0.389.2$\pm$ 0.389.2 ± 0.3 / 2.3×2.3$\times$2.3 × 90.7±0.290.7$\pm$ 0.290.7 ± 0.2 / 2.9×2.9$\times$2.9 ×
speaker-id 83.2±0.283.2$\pm$ 0.283.2 ± 0.2 / 1×1$\times$1 × 79.6±0.579.6$\pm$ 0.579.6 ± 0.5 / 2.9×2.9$\times$2.9 × 82.4±0.882.4$\pm$ 0.882.4 ± 0.8 / 3.7×3.7$\times$3.7 ×
Average gain 1×1$\times$1 × 4.2×4.2$\times$4.2 × 4.5×4.5$\times$4.5 ×
5.4 Ablation studies